A car travelling at 60km/h overtakes another car travelling at 42km/h.Assuming each car to be 5m long, find the time taken during the overtake and the total road distance used for the overtake.

Grade:12

1 Answers

Pratham Ashish
14 years ago

hi ankur,

1st of all calculate the relative velocity between these 2 cars ,ie, (60-42) = 18 km/hr.

now mke unit conversion then relative velocity becomes 5m/s.......

since to overtake the 2nd car has to travel 10m as 5m is the length of each car , so 5m to catch up the other car and 5m to clear it.

 

time taken to do so ,t = 10/5 = 2 sec

to find road distance ,we simply calculate distance travelled by 2nd car...

s = ut

u = 60km/h = 16.7m/s

s = 16.7 *2 = 33.3m